In a right triangle, the length of the longer leg is 7 more inches than the shorter leg. The length of the hypotenuse is 8 inches more than the shorter leg. Find the length of the hypotenuse.
Answer by Cromlix(4381)

Hi there,
Short side = x
Long side = 7 + x
Hypotenuse = 8 + x
By Pythagoras Theorem
x^2 + (7 + x)^2 = (8 + x)^2
x^2 + 49 + 14x + x^2 = 64 + 16x + x^2
x^2 - 2x - 15 = 0
(x + 3)(x - 5) = 0
x + 3 = o
x = -3 (Disregard as -ve)
x - 5 = 0
x = 5
Short side = 5 inches
Long side = 12 inches
Hypotenuse = 13 inches
Hope this helps :-)